A rocket of mass 0.41 kg is set on the ground, pointing vertically up.
When ignited, the gunpowder burns for a period of 1.6 seconds, and provides a constant force of 16 N.
The mass of the gunpowder is very small compared to the mass of the rocket, and can be ignored.
The effects of air resistance can also be ignored. Use g = 9.8 m s-2. Answer accurate to 0.1 N.

What is the magnitude of the resultant force on the rocket?
1
Expert's answer
2020-01-27T09:23:16-0500

Mass of the rocket=0.41kg

Force=16N

Let the final acceleration of the rocket is "a m\/sec^2"

Now, "ma=F-mg"

"0.41a=16-0.41\times 9.8"

"\Rightarrow F_{net}=11.98N =12.0N"